OpenStudy (annie96):

Select the equation of the line in the slope-intercept form: 4x+9y=36 y=-4/9x-4 y=-4/9x+4 y=-9/4x-4 y=9/4x+4

OpenStudy (anonymous):

4x+9y=36 Bring all terms to the right hand side , leaving y term on the left hand side. 9y = 36-4x Make y into one thorughout y=4-4/9x y= -4/9x+4

OpenStudy (annie96):

thank you.